Find the quotient and the remainder of the function
F(x) =
3x4 + 8x3 + 7x + 2
x + 2
3
i) long division (5marks)
ii) Synthetic division (3marks)
b. i) Factorize completely F(x) = x3 + 9x2 + 23x + 15 (4marks)
ii) Show if the F(x) = 3x4 +8x3 +7x2
1
Expert's answer
2020-12-11T12:39:06-0500

i) long division


"\begin{matrix}
 & \underline{3x^3+2x^2-4x+15} \\
 x+2&)3x^4+8x^3+7x+2\ \ \ \\
-
\end{matrix}""\begin{matrix}
 & \underline{3x^4+6x^3 }&&\ \ \\
 & \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 2x^3+0x^2+7x+2 \\
 
\end{matrix}""-\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \""\begin{matrix}
 & \underline{2x^3+4x^2 }&& \\
 & \ \ \ -4x^2+7x+2 
\end{matrix}""-\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \""\begin{matrix}
 & \underline{-4x^2-8x}&&& \\
 & \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 15x+2 
\end{matrix}""-\ \ \ \ \""\begin{matrix}
 &\ \ \ \ \ \ \ \ \ \ \ \ \underline{15x+30} \\
 & \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ -28 
\end{matrix}"

"\dfrac{3x^4+8x^3+7x+2}{x+2}=3x^3+2x^2-4x+15+\dfrac{-28}{x+2}"


"Quotient=3x^3+2x^2-4x+15"

"Remainder=-28"


ii) Synthetic division


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 & x^4 & x^3 & x^2 & x^1 & x^0 \\ \hline
 -2 & 3 & 8 & 0 & 7 & 2 \\
 & & -6 & -4 & 8 & -30 \\
 \hdashline
 & 3 & 2 & -4 &15 & -28
\end{array}"

"\dfrac{3x^4+8x^3+7x+2}{x+2}=3x^3+2x^2-4x+15+\dfrac{-28}{x+2}"



"Quotient=3x^3+2x^2-4x+15"

"Remainder=-28"

b) i)

Synthetic division


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 & x^3 & x^2 & x^1 & x^0 \\ \hline
 -1 & 1 & 9 & 23 & 15 \\
 & & -1 & -8 & 8 & \\
 \hdashline
 & 1 & 8 & 15 &0
\end{array}"

"\dfrac{x^3+9x^2+23x+15}{x+1}=x^2+8x+15"



"Quotient=x^2+8x+15"

"Remainder=0"


"x^2+8x+15=(x+5)(x+3)"



"x^3+9x^2+23x+15=(x+5)(x+3)(x+1)"

ii) Show if the F(x) = 3x4 +8x3 +7x2

Synthetic division


"\def\arraystretch{1.5}
 \begin{array}{c:c}
 & x^4 & x^3 & x^2 & x^1 & x^0 \\ \hline
 -2 & 3 & 8 & 7 & 0 & 0 \\
 & & -6 & -4 & -6 & 12 \\
 \hdashline
 & 3 & 2 & 3 &-6 & 12
\end{array}"

"\dfrac{3x^4+8x^3+7x^2}{x+2}=3x^3+2x^2+3x-6+\dfrac{12}{x+2}"



"Quotient=3x^3+2x^2+3x-6"

"Remainder=12"